# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 3.2 The Clebsch-Gordan decomposition
$C_{1}e^{-\bm{Y}_{1}+2\bm{X}_{1}}C_{1}^{-1}$
$=$ $\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)^{-1}e^{\bm{X}_{1}}e^{-\bm{Y}_{1}}e^{\bm{X}_{1}}\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}$
$=$ $e^{\bm{X}_{1}}\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\varphi\left({\bf q}^{-1}e^{\bm{Y}_{1}-L_{2}}\right)^{-1}e^{-\bm{Y}_{1}}\varphi\left({\bf q}e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}e^{\bm{X}_{1}}$
$=$ $e^{\bm{X}_{1}}\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}e^{-\bm{Y}_{1}}\left(1+e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}e^{\bm{X}_{1}}$
$=$ $e^{\bm{X}_{1}}\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\left(e^{-\bm{Y}_{1}}+e^{-L_{2}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}e^{\bm{X}_{1}}$
$=$ $\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}e^{\bm{X}_{1}}e^{-\bm{Y}_{1}}e^{\bm{X}_{1}}\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}+e^{2\bm{X}_{1}-L_{2}}.$
@@
$C_{1}e^{-\bm{Y}_{1}}Q_{1}C_{1}^{-1}=$ $\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}e^{-\bm{Y}_{1}}\left(e^{L_{1}}+e^{-L_{1}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}.$
@@
$=$ $e^{-2\bm{X}_{1}+L_{2}}+e^{2\bm{X}_{1}-L_{2}}+\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\left({\bf q}^{-1}e^{-2\bm{X}_{1}}e^{-\bm{Y}_{1}}+{\bf q}e^{2\bm{X}_{1}}e^{-\bm{Y}_{1}}+e^{-\bm{Y}_{1}+L_{1}}+e^{-\bm{Y}_{1}-L_{1}}\right)\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}$
$=$ $e^{-2\bm{X}_{1}+L_{2}}+e^{2\bm{X}_{1}-L_{2}}+\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\left({\bf q}^{-1}e^{-2\bm{X}_{1}+L_{1}}+{\bf q}e^{2\bm{X}_{1}+L_{1}}+e^{2L_{1}}+1\right)e^{-\bm{Y}_{1}-L_{1}}\frac{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}$
$=$ $e^{-2\bm{X}_{1}+L_{2}}+e^{2\bm{X}_{1}-L_{2}}+\frac{\varphi\left(e^{2\bm{X}_{1}+L_{1}}\right)}{\varphi\left(e^{-2\bm{X}_{1}+L_{1}}\right)}\left(1+{\bf q}e^{2\bm{X}_{1}+L_{1}}\right)\left(1+{\bf q}^{-1}e^{-2\bm{X}_{1}+L_{1}}\right)\frac{\varphi\left({\bf q}^{-2}e^{-2\bm{X}_{1}+L_{1}}\right)}{\varphi\left({\bf q}^{2}e^{2\bm{X}_{1}+L_{1}}\right)}e^{-\bm{Y}_{1}-L_{1}}$
$=$ $e^{-2\bm{X}_{1}+L_{2}}+e^{2\bm{X}_{1}-L_{2}}+e^{-\bm{Y}_{1}-L_{1}}.$
@@
$Q_{1}^{\prime\prime}=\mathcal{U}_{21}Q_{21}\mathcal{U}_{21}^{-1}=e^{-2\bm{X}_{1}+L_{2}}+e^{2\bm{X}_{1}-L_{2}}+e^{-\bm{Y}_{1}-L_{1}}.$
A similar computation gives for the tilded operator
$\tilde{Q}_{1}^{\prime\prime}=\mathcal{U}_{21}\tilde{Q}_{21}\mathcal{U}_{21}^{-1}=e^{-2\tilde{\bm{X}}_{1}+\tilde{L}_{2}}+e^{2\tilde{\bm{X}}_{1}-\tilde{L}_{2}}+e^{-\tilde{\bm{Y}}_{1}-\tilde{L}_{1}}.$
@@
If $\dim\mathcal{H}\left(\ell,\ell^{*}\right)=1$, each $\bm{\mathcal{H}}\left(\ell,\ell^{*}\right)$ is a irreducible representation of $(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}})\circ\Delta$
@@
$F_{2},E_{2},K_{2},K_{2}^{-1}$ and their tilded relatives are represented irreducibly on the first factor in $\bm{\mathcal{H}}\left(\ell,\ell^{*}\right)=\mathcal{H}\otimes\mathcal{H}\left(\ell,\ell^{*}\right)$, and they are explicitly given by
$K_{2}$ $=e^{-\bm{X}_{2}},\qquad K_{\lambda}^{-1}=e^{\bm{X}_{2}},\qquad E_{2}=-\frac{1}{{\bf q}-{\bf q}^{-1}}e^{-\bm{X}_{2}+\bm{Y}_{2}},$
$F_{2}$ $=\frac{1}{\left({\bf q}-{\bf q}^{-1}\right)^{2}}\left({\bf q}^{-1}K_{2}^{2}+{\bf q}K_{2}^{-2}+Q_{1}^{\prime\prime}\right)E_{2}^{-1},\qquad Q_{1}^{\prime\prime}=\ell\mathrm{id}_{\mathcal{H}}$
@@
$\mathcal{U}_{21}^{-1}K_{2}\mathcal{U}_{21}$ $=t_{21}\mathcal{S}_{2}e^{-\bm{X}_{2}}\mathcal{S}_{2}^{-1}t_{21}^{-1}=t_{21}e^{-\bm{X}_{2}-\bm{X}_{1}}t_{21}^{-1}$
$=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-\bm{X}_{2}-\bm{X}_{1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}$
$=e^{-\bm{X}_{2}-\bm{X}_{1}}=K_{2}K_{1}=\left(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}}\right)\circ\Delta(K)$
$\mathcal{U}_{21}^{-1}E_{2}\mathcal{U}_{21}$ $=-\frac{1}{{\bf q}-{\bf q}^{-1}}t_{21}\mathcal{S}_{2}e^{-\bm{X}_{2}+\bm{Y}_{2}}\mathcal{S}_{2}^{-1}t_{21}^{-1}$
$=-\frac{1}{{\bf q}-{\bf q}^{-1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-\bm{X}_{1}-\bm{X}_{2}+\bm{Y}_{2}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}$
$=-\frac{1}{{\bf q}-{\bf q}^{-1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)\varphi\left({\bf q}^{-2}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)^{-1}e^{-\bm{X}_{1}-\bm{X}_{2}+\bm{Y}_{2}}$
$=-\frac{1}{{\bf q}-{\bf q}^{-1}}\left(1+{\bf q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+2\bm{X}_{2}}\right)e^{-\bm{X}_{1}-\bm{X}_{2}+\bm{Y}_{2}}$
$=-\frac{1}{{\bf q}-{\bf q}^{-1}}\left(e^{-\bm{X}_{1}-\bm{X}_{2}+\bm{Y}_{2}}+e^{\bm{Y}_{1}-\bm{X}_{1}+\bm{X}_{2}}\right)=E_{2}K_{1}+K_{2}^{-1}E_{1}$
$=\left(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}}\right)\circ\Delta(E)$
Combining the above result $\mathcal{U}_{21}^{-1}Q_{1}^{\prime\prime}\mathcal{U}_{21}=Q_{21}=(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}})\circ\Delta Q$,
$\left(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}}\right)\circ\Delta(F)$ $=\left(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}}\right)\circ\Delta\frac{\left({\bf q}^{-1}K^{2}+{\bf q}K^{-2}+Q\right)E^{-1}}{\left({\bf q}-{\bf q}^{-1}\right)^{2}}$
$=\mathcal{U}_{21}^{-1}\frac{\left({\bf q}^{-1}K_{2}^{2}+{\bf q}K_{2}^{-2}+Q_{1}^{\prime\prime}\right)E_{2}^{-1}}{\left({\bf q}-{\bf q}^{-1}\right)^{2}}\mathcal{U}_{21}=\mathcal{U}_{21}^{-1}F_{2}\mathcal{U}_{21}.$
@@
The above shows that if If $\dim\mathcal{H}\left(\ell,\ell^{*}\right)=1$ holds, the unitary transformation $\mathcal{U}_{21}$ is a Clebsch-Gordan map intertwining three representations $\lambda_{1},\lambda_{2},\lambda_{3}$:
@@
$\mathcal{U}_{21}\left[\left(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}}\right)\circ\Delta(X)\right]\mathcal{U}_{21}^{-1}=\pi_{\lambda_{3},\tilde{\lambda}_{3}}(X)$ (3.22)
where $\lambda_{3}+\lambda_{3}^{-1}=\ell$ and $\tilde{\lambda}_{3}+\tilde{\lambda}_{3}^{-1}=\ell^{*}$.
The property $\dim\mathcal{H}\left(\ell,\ell^{*}\right)=1$ can be derived if we assume $\lambda_{a}=e^{L_{a}}$, $a=1,2$, is quantized
$\lambda_{a}=\exp\left[\frac{2\pi i}{k}(-ib\mu_{L_{a}}-m_{L_{a}})\right],\qquad\mu_{L_{a}}\in\mathbb{R},\ m_{L_{a}}\in\mathbb{Z}/k\mathbb{Z}.$ (3.23)
Then we can further apply a unitary shift operator
$\mathcal{S}_{L_{1}}=\exp\left[\frac{2\pi i}{k}\left(\mu_{L_{1}}\bm{\nu}_{1}-m_{L_{1}}\bm{n}_{1}\right)\right],\qquad\mathcal{S}_{L_{1}}e^{-\bm{Y}_{1}-L_{1}}\mathcal{S}_{L_{1}}^{-1}=e^{-\bm{Y}_{1}}.$ (3.24)
We obtain
$\mathcal{Q}_{1}^{\prime\prime}=\mathcal{S}_{L_{1}}Q_{1}^{\prime\prime}\mathcal{S}_{L_{1}}^{-1}$ $=$ $e^{-2\bm{X}_{1}+L_{2}}+e^{2\bm{X}_{1}-L_{2}}+e^{-\bm{Y}_{1}},$ (3.25)
$\tilde{\mathcal{Q}}_{1}^{\prime\prime}=\mathcal{S}_{L_{1}}\tilde{Q}_{1}^{\prime\prime}\mathcal{S}_{L_{1}}^{-1}$ $=$ $e^{-2\tilde{\bm{X}}_{1}+\tilde{L}_{2}}+e^{2\tilde{\bm{X}}_{1}-\tilde{L}_{2}}+e^{-\tilde{\bm{Y}}_{1}}.$ (3.26)
@@
The spectral decomposition of $\mathcal{Q}_{1}^{\prime\prime},\tilde{\mathcal{Q}}_{1}^{\prime\prime}$ will be discussed in detail in Section 5. The analysis shows that $\dim\mathcal{H}(\ell,\ell^{*})=1$ indeed holds.
4 Quantum flat connections on 4-holed sphere

### v2

@@ 3.2 The Clebsch-Gordan decomposition
$C_{1}e^{-\bm{Y}_{1}+\bm{U}_{1}}C_{1}^{-1}$
$=$ $\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\varphi\left(e^{\bm{Y}_{1}-L_{2}}\right)^{-1}{\bf q}^{-1}e^{-\bm{Y}_{1}}\varphi\left({\bf q}^{2}e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}e^{\bm{U}_{1}}$
$=$ $\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\bf q}^{-1}e^{-\bm{Y}_{1}}\left(1+{\bf q}e^{\bm{Y}_{1}-L_{2}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}e^{\bm{U}_{1}}$
$=$ $\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\left({\bf q}e^{\bm{U}_{1}}e^{-\bm{Y}_{1}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}+e^{\bm{U}_{1}-L_{2}}$
@@
$C_{1}e^{-\bm{Y}_{1}}Q_{1}C_{1}^{-1}=$ $\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}e^{-\bm{Y}_{1}}\left(e^{L_{1}}+e^{-L_{1}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}.$
@@
$=$ $e^{-\bm{U}_{1}+L_{2}}+e^{\bm{U}_{1}-L_{2}}+\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\left({\bf q}^{-1}e^{-\bm{U}_{1}}e^{-\bm{Y}_{1}}+{\bf q}e^{\bm{U}_{1}}e^{-\bm{Y}_{1}}+e^{-\bm{Y}_{1}+L_{1}}+e^{-\bm{Y}_{1}-L_{1}}\right)\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}$
$=$ $e^{-\bm{U}_{1}+L_{2}}+e^{\bm{U}_{1}-L_{2}}+\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\left({\bf q}^{-1}e^{-\bm{U}_{1}+L_{1}}+{\bf q}e^{\bm{U}_{1}+L_{1}}+e^{2L_{1}}+1\right)e^{-\bm{Y}_{1}-L_{1}}\frac{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}$
$=$ $e^{-\bm{U}_{1}+L_{2}}+e^{\bm{U}_{1}-L_{2}}+\frac{\varphi\left(e^{\bm{U}_{1}+L_{1}}\right)}{\varphi\left(e^{-\bm{U}_{1}+L_{1}}\right)}\left(1+{\bf q}e^{\bm{U}_{1}+L_{1}}\right)\left(1+{\bf q}^{-1}e^{-\bm{U}_{1}+L_{1}}\right)\frac{\varphi\left({\bf q}^{-2}e^{-\bm{U}_{1}+L_{1}}\right)}{\varphi\left({\bf q}^{2}e^{\bm{U}_{1}+L_{1}}\right)}e^{-\bm{Y}_{1}-L_{1}}$
$=$ $e^{-\bm{U}_{1}+L_{2}}+e^{\bm{U}_{1}-L_{2}}+e^{-\bm{Y}_{1}-L_{1}}.$
@@
$Q_{1}^{\prime\prime}=\mathcal{U}_{21}Q_{21}\mathcal{U}_{21}^{-1}=e^{-\bm{U}_{1}+L_{2}}+e^{\bm{U}_{1}-L_{2}}+e^{-\bm{Y}_{1}-L_{1}}.$
The Hermitian conjugate gives for the tilded operator
$\tilde{Q}_{1}^{\prime\prime}=\mathcal{U}_{21}\tilde{Q}_{21}\mathcal{U}_{21}^{-1}=e^{-\tilde{\bm{U}}_{1}+\tilde{L}_{2}}+e^{\tilde{\bm{U}}_{1}-\tilde{L}_{2}}+e^{-\tilde{\bm{Y}}_{1}-\tilde{L}_{1}}.$
$Q_{1}^{\prime\prime}$ and $\tilde{Q}_{1}^{\prime\prime}$ commute, so they are normal operators. $Q_{1}^{\prime\prime}$ and $\tilde{Q}_{1}^{\prime\prime}$ only acting on the second copy of $\mathcal{H}\otimes\mathcal{H}$, and their spectral decomposition gives a direct integral decomposition
@@
where $\mu$ is the spectral measure, $P_{\lambda,\tilde{\lambda}}$ is the projective-valued measure, and $\tilde{\lambda}=\lambda^{*}$. Therefore,
@@
[…]
@@
[…]
$K_{2}$ $=e^{-\bm{U}_{2}},\qquad K_{\lambda}^{-1}=e^{\bm{U}_{2}},\qquad F_{2}=-\frac{i}{{\bf q}-{\bf q}^{-1}}qe^{\bm{Y}_{2}},$
$E_{2}$ $=-\frac{1}{\left({\bf q}-{\bf q}^{-1}\right)^{2}}\left({\bf q}^{-1}K_{2}+{\bf q}K_{2}^{-1}+Q_{1}^{\prime\prime}\right)F_{2}^{-1},\qquad Q_{1}^{\prime\prime}=(\lambda+\lambda^{-1})\mathrm{id}_{\mathcal{H}}$
@@
$\mathcal{U}_{21}^{-1}K_{2}\mathcal{U}_{21}$ $=t_{21}\mathcal{S}_{2}e^{-\bm{U}_{2}}\mathcal{S}_{2}^{-1}t_{21}^{-1}=t_{21}e^{-\bm{U}_{2}-\bm{U}_{1}}t_{21}^{-1}$
$=\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{-\bm{U}_{2}-\bm{U}_{1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}$
$=e^{-\bm{U}_{2}-\bm{U}_{1}}=K_{2}K_{1}=\left(\Delta K\right)_{21}$
$\mathcal{U}_{21}^{-1}F_{2}\mathcal{U}_{21}$ $=-\frac{iq}{{\bf q}-{\bf q}^{-1}}t_{21}\mathcal{S}_{2}e^{\bm{Y}_{2}}\mathcal{S}_{2}^{-1}t_{21}^{-1}$
$=-\frac{iq}{{\bf q}-{\bf q}^{-1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{\bm{Y}_{2}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}$
$=-\frac{iq}{{\bf q}-{\bf q}^{-1}}\varphi\left(e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)\varphi\left(\mathbf{q}^{-2}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)^{-1}e^{\bm{Y}_{2}}$
$=-\frac{iq}{{\bf q}-{\bf q}^{-1}}\left(1+\mathbf{q}^{-1}e^{\bm{Y}_{1}-\bm{Y}_{2}+\bm{U}_{2}}\right)e^{\bm{Y}_{2}}$
$=-\frac{iq}{{\bf q}-{\bf q}^{-1}}\left(e^{\bm{Y}_{2}}+e^{\bm{Y}_{1}+\bm{U}_{2}}\right)=F_{2}+K_{2}^{-1}F_{1}$
$=\left(\Delta F\right)_{21}$
Moreover we have $\mathcal{U}_{21}^{-1}F_{2}^{-1}\mathcal{U}_{12}=(\Delta F)_{21}^{-1}$. Therefore, at the level of the representation, $Q_{21}=-\left({\bf q}-{\bf q}^{-1}\right)^{2}(\Delta E)_{21}(\Delta F)_{21}-\mathbf{q}^{-1}(\Delta K)_{21}-{\bf q}(\Delta K)_{21}^{-1}$ implies
$(\Delta E)_{21}$ $=$ $-\frac{\left[\mathbf{q}^{-1}(\Delta K)_{21}+{\bf q}(\Delta K)_{21}^{-1}+Q_{21}\right](\Delta F)_{21}^{-1}}{\left({\bf q}-{\bf q}^{-1}\right)^{2}}$ (3.35)
$=$ $\mathcal{U}_{21}^{-1}\frac{\left({\bf q}^{-1}K_{2}+{\bf q}K_{2}^{-1}+Q_{1}^{\prime\prime}\right)F_{2}^{-1}}{\left({\bf q}-{\bf q}^{-1}\right)^{2}}\mathcal{U}_{21}=\mathcal{U}_{21}^{-1}E_{2}\mathcal{U}_{21},$
by using $\mathcal{U}_{21}^{-1}Q_{1}^{\prime\prime}\mathcal{U}_{21}=Q_{21}$. ∎
@@
When $\dim\mathcal{H}\left(\ell,\ell^{*}\right)=1$ holds, the unitary transformation $\mathcal{U}_{21}$ is a Clebsch-Gordan map intertwining three representations $\lambda_{1},\lambda_{2},\lambda_{3}$:
@@
$\mathcal{U}_{21}\left[\left(\pi_{\lambda_{2},\tilde{\lambda}_{2}}\otimes\pi_{\lambda_{1},\tilde{\lambda}_{1}}\right)\circ\Delta(X)\right]\mathcal{U}_{21}^{-1}=\int_{\mathbb{C}}^{\oplus}\pi_{\lambda_{3},\tilde{\lambda}_{3}}(X)\otimes\mathrm{d}P_{\lambda_{3},\tilde{\lambda}_{3}}.$ (3.36)
In the following, we assume $\lambda_{a}=e^{L_{a}}$, $a=1,2$, is quantized
$\lambda_{a}=\exp\left[\frac{2\pi i}{N}(-ib\mu_{a}-m_{a})\right],\qquad\mu_{a}\in\mathbb{R},\ m_{a}\in\mathbb{Z}/N\mathbb{Z}.$ (3.37)
We define the following unitary transformations
$\mathcal{S}_{\lambda_{1}}f\left(\mu,m\right)=f\left(\mu-\mu_{1},m-m_{1}\right),\qquad{\mathcal{D}}_{\lambda_{2}}f\left(\mu,m\right)=e^{-\frac{2\pi i}{N}(\mu_{2}\mu-m_{2}m)}f\left(\mu,m\right),$ (3.38)
satisfying $\mathcal{S}_{\lambda_{1}}\mathcal{D}_{\lambda_{2}}=e^{\frac{2\pi i}{N}(\mu_{1}\mu_{2}-m_{1}m_{2})}\mathcal{D}_{\lambda_{2}}\mathcal{S}_{\lambda_{1}}$ and
$\mathcal{S}_{\lambda_{1}}\bm{y}^{-1}\mathcal{S}_{\lambda_{1}}^{-1}=\lambda_{1}\bm{y}^{-1},\quad\mathcal{S}_{\lambda_{1}}\bm{u}\mathcal{S}_{\lambda_{1}}^{-1}=\bm{u},\quad\mathcal{S}_{\lambda_{1}}\tilde{\bm{y}}^{-1}\mathcal{S}_{\lambda_{1}}^{-1}={\lambda}^{*}_{1}\tilde{\bm{y}}^{-1},\quad\mathcal{S}_{\lambda_{1}}\tilde{\bm{u}}\mathcal{S}_{\lambda_{1}}^{-1}=\tilde{\bm{u}},$ (3.39)
$\mathcal{D}_{\lambda_{2}}^{-1}\bm{y}\mathcal{D}_{\lambda_{2}}=\bm{y},\qquad\mathcal{D}_{\lambda_{2}}^{-1}\bm{u}\mathcal{D}_{\lambda_{2}}=\lambda_{2}\bm{u},\qquad\mathcal{D}_{\lambda_{2}}^{-1}\tilde{\bm{y}}\mathcal{D}_{\lambda_{2}}=\tilde{\bm{y}},\qquad\mathcal{D}_{\lambda_{2}}^{-1}\tilde{\bm{u}}\mathcal{D}_{\lambda_{2}}={\lambda}^{*}_{2}\tilde{\bm{u}}.$ (3.40)
The Fourier transformation
[…]
satisfies
$\mathcal{F}\bm{u}\mathcal{F}^{-1}=\bm{y},\quad\mathcal{F}\bm{y}^{-1}\mathcal{F}^{-1}=\bm{u},\quad\mathcal{F}\tilde{\bm{u}}\mathcal{F}^{-1}=\tilde{\bm{y}},\quad\mathcal{F}\tilde{\bm{y}}^{-1}\mathcal{F}^{-1}=\tilde{\bm{u}}.$ (3.42)
Applying these unitary transformations to $Q^{\prime\prime}_{2}$ and $\tilde{Q}_{2}^{\prime\prime}$ leads to the simplification
$\bm{L}\equiv{\cal F}\mathcal{D}_{\lambda_{2}}^{-1}\mathcal{S}_{\lambda_{1}}Q_{2}^{\prime\prime}\mathcal{S}_{\lambda_{1}}^{-1}\mathcal{D}_{\lambda_{2}}{\cal F}^{-1}$ $=$ $\bm{y}^{-1}+\bm{y}+\bm{u},$ (3.43)
$\tilde{\bm{L}}\equiv{\cal F}\mathcal{D}_{\lambda_{2}}^{-1}\mathcal{S}_{\lambda_{1}}\tilde{Q}_{2}^{\prime\prime}\mathcal{S}_{\lambda_{1}}^{-1}\mathcal{D}_{\lambda_{2}}{\cal F}^{-1}$ $=$ $\tilde{\bm{y}}^{-1}+\tilde{\bm{y}}+\tilde{\bm{u}}.$ (3.44)
These pair of operators are generalizations to the Dehn-twist operators in quantum Teichmüller theory, which corresponds to $N=1$ [1, 7]. The Dehn-twist operators are unitary equivalent to the Casimir operators of the tensor product representations for the modular double of $U_{q}(sl(2,\mathbb{R}))$ [4, 7, 3].
@@
The spectral decomposition of $\bm{L},\tilde{\bm{L}}$ will be discussed in detail in Section 5. The analysis shows that $\dim\mathcal{H}(\ell,\ell^{*})=1$ indeed holds.
4 Quantization of flat connections on 4-holed sphere